OpenStudy (anonymous):

Solve the equation 4 + 6c = 6 - 4c

OpenStudy (saifoo.khan):

6c + 4c = 6 -4 10c = 2 c = 2/10 c = 1/5

OpenStudy (anonymous):

10c = 2 c = 1/5